OpenStudy (anonymous):

Find the equation of the line in slope-intercept form (y= mx+b) M=3/2 and passes through (-6,2)

OpenStudy (tkhunny):

Have you considered substituting the known values and solving for whatever is left?

OpenStudy (anonymous):

No i actially dont know this its confusing

OpenStudy (liv1234):

Do you know how to use the slope intercept form?

OpenStudy (anonymous):

Yes but not for this

OpenStudy (liv1234):

Okay, I can help you. Just give me a minute.

OpenStudy (anonymous):

Thanks❤️ Math is not my bestfriend

OpenStudy (liv1234):

Y=3/2x+2 I believe that's how you put it into the equation.

OpenStudy (anonymous):

Ugh i gotta show my work on how i got it

OpenStudy (liv1234):

You have your numbers already. M=3/2 (-6,2)

OpenStudy (tkhunny):

Seriously, just substitute. Here's your equation. y = mx+b You have m = 3/2 -- Substitute This gives y = (3/2)x + b You have (-6,2) ==> x = -6 and y = 2 -- substitute 2 = (3/2)(-6) + b Can you solve for b?

OpenStudy (liv1234):

\[(3/2)*(-6)=?\]

OpenStudy (tkhunny):

You must DO some work in order to show it. Let's see what you can do.

OpenStudy (liv1234):

I was just writing that out to make it a little easier for them.

OpenStudy (anonymous):

Thanks i can solve for B

OpenStudy (anonymous):

Thanksss so much

OpenStudy (liv1234):

What did you get after multiplying?

OpenStudy (anonymous):

Im doing it now

OpenStudy (liv1234):

Alright. Let me know.

OpenStudy (liv1234):

You should have gotten -9.

OpenStudy (anonymous):

Imma show you how my teacherr did it and YESS YAAY i got that answer

OpenStudy (liv1234):

2 = (3/2)(-6) + b b = -9 2 = (3/2)(-6)+-9

OpenStudy (liv1234):

After doing that you should get -18. 2=-18 Divide each side by 2. 2/2=1 -18/2=-9

OpenStudy (tkhunny):

Well, if your teacher gave you y = mx+b and then gave you M = 3/2 We have a problem. Changing from upper case to lower case in the middle of a problem just is not good form.

OpenStudy (anonymous):

Heres some examples my teacher did .... Its confusing

OpenStudy (anonymous):

See

OpenStudy (tkhunny):

These are all very similar. You just have to recognize the information you are given. In the first, you are just given the slope. m = something In the second, you are just given the slope. It's in disguise. Parallel mean SAME slope. Find the slope of the given line and use that for m. In the third, you are just given the slope. It's in disguise. Perpendicular mean the slope of this line, multiplied by the slope of the desired line, give -1. Find the slope of the given line and calculate the slope that you need. Examples. If the slope of the perpendicular line is 2, then the slope of your line is -1/2, since 2 * (-1/2) = -1 If the slope of the perpendicular line is 1/3, then the slope of your line is -3, since (1/3) * (-3) = -1 Post the next one (parallel slopes) on a new thread and show your best work.